OpenStudy (anonymous):

using the quadractic formula solve 3x^2-17=0

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

hartnn (hartnn):

can u find a,b,c ?

OpenStudy (anonymous):

no

hartnn (hartnn):

a=3,b=0,c=-17 now just put these in the formula i gave ....

OpenStudy (anonymous):

oh ok i get it now

hartnn (hartnn):

good, so what 2 values of x u got ?

OpenStudy (anonymous):

i got 2.38 and -2.38

hartnn (hartnn):

that is correct! good work :)

OpenStudy (anonymous):

how about this equation 2y^2=-5y

hartnn (hartnn):

here a=2, b= 5,c=0 do u get how ?

OpenStudy (anonymous):

so would the answers be 2.28 and .22

OpenStudy (anonymous):

in the back of the book they have -5/2,0

hartnn (hartnn):

do u need to use formula ? because there is simpler way for 2y^2=-5y

OpenStudy (anonymous):

in the book it says that you do

hartnn (hartnn):

so did u get how a=2, b= 5,c=0

hartnn (hartnn):

then u just put these in formula

OpenStudy (anonymous):

i did put them in the formula but how do you get -5/2 and 0?

hartnn (hartnn):

ok, what u got b^2-4ac as ?

hartnn (hartnn):

5^2-4*2*0 = ?

OpenStudy (anonymous):

yes i have that

hartnn (hartnn):

so what numerator u got ?

hartnn (hartnn):

the denominator =2a=4

hartnn (hartnn):

\(\large -b+\sqrt{b^2-4ac}=-5+\sqrt{25-0}=-5+5=0\) did u get this ?

hartnn (hartnn):

so one of the root = 0

hartnn (hartnn):

\(\large -b-\sqrt{b^2-4ac}=-5-\sqrt{25-0}=-5-5=-10\) and with denominator =4, u get other root as -10/1 = -2/5 got this ?

hartnn (hartnn):

*-10/4=-5/2

OpenStudy (anonymous):

ok

hartnn (hartnn):

good, any more doubts ?

OpenStudy (anonymous):

no that is it

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^